OpenStudy (anonymous):

Solve, using linear combination. 3x + y = 4 2x + y = 4 A. (–1, 7) B. (0, 4) C. (–3, 12) D. (–3, 13)

OpenStudy (anonymous):

3x + y = 4 -2x -y = -4 (multipled by -1) +3x + y = 4 -2x -y = -4 x + 0 = 0 x=0

OpenStudy (anonymous):

and now put for x= 0 3(0)+y=4 and you get y=4